# HERA: High-efficiency Matrix Compression via Element Replacement

### v2

@@ Abstract
We propose HETA (HETA) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix. This matrix is then grouped and quantized by columns. To enhance HETA, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
Experimental results demonstrate that HETA can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms. Our contributions include the abstraction of the QEM problem, the design of the HETA algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@ 1 Introduction
There has been a growing literature on the matrix quantization. The first category of methods focuses on independently compressing the elements of a matrix for each specific scenario. Examples of such techniques include LLM.int8 (Ge et al. (2013)), Optimal Brain Damage (LeCun et al. (1989)), GPTQ (Frantar et al. (2022)), AWQ (Lin et al. (2023)), SmoothQuant (Xiao et al. (2023)), and OmniQuant (Shao et al. (2023)). The second category groups matrix elements by columns and then applies quantization to each group. Relevant works in this category include product quantization (PQ) (Jegou et al. (2010)), optimized product quantization (OPQ) (Ge et al. (2013)), locally optimized product quantization (LOPQ) (Kalantidis & Avrithis (2014)), and SqueezeLLM (Kim et al. (2023)). The first category of work can be summarized as utilizing outliers and the importance of elements in specific scenarios to improve the RTN (Round-To-Nearest)  [^1] RTN (Round-To-Nearest) is an algorithm that quantizes elements by rounding each value to its nearest representable level. (Gray & Neuhoff (1998) algorithm). The second category focuses on enhancements to PQ algorithms. The primary drawback of the first category is that it does not consider the correlation between elements, instead independently quantizing and storing elements. Conversely, the second category fails to account for the relative order of elements.
@@
Using these ideas, we propose HETA (HETA) to address the QEM problem in matrix quantization. The core idea of HETA is to leverage the local orderliness of matrix elements to optimize the QEM problem. The design involves swapping adjacent elements of the matrix to form a new, locally ordered matrix. To cover a broader range, we can perform further element swapping on the initially locally ordered matrix, similar to the approach of receptive field used in convolutional neural networks (LeCun et al. (1998)). This step can be repeated for multiple iterations. The newly ordered matrix is then grouped by columns, with each group being quantized.
Additionally, we propose two optimization algorithms based on the basic HETA. First, the residuals between the original and quantized matrices can be further quantized to enhance the accuracy of the results. Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the HETA.
We evaluate HETA on multiple datasets. Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
@@
  • Designed an algorithm: We developed the HETA (HETA) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate HETA algorithm.
@@ 3 METHODS
The core idea of HETA (HETA) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm. Additionally, multiple iterations are used to expand the algorithm’s swapping field. We begin by describing the basic version of this idea and then propose two optimizations for the basic algorithm.
@@ 3.1 Basic algorithm
The steps and design of the HETA algorithm are illustrated in Figure 1. The figure provides a conceptual overview of the HETA (HETA) algorithm.
On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the HETA is depicted. This iterative procedure involves comparing adjacent elements and ensuring that adjacent elements are not placed in the same matrix. This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”HETA.” The HETA algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values. This approach is based on two observations. First, the local orderliness of the matrix can enhance the regularity of matrix elements, thereby improving compression efficiency. Second, multiple iterations allow the orderliness to cover a larger number of elements. For instance, adjacent elements after the first iteration might be non-adjacent in the original matrix. This iterative process, similar to that of convolutional neural networks, extends the ”receptive field” to cover more elements.
On the right side of the figure, the HETA compression process is shown. After the elements have been rearranged, the algorithm groups the sub-matrices, followed by clustering to determine the centroids. These centroids are used to quantize the matrix. The final quantized matrix, codebook, and indicator maps are the outputs of the HETA algorithm. This divide-and-conquer compression method is based on an observation: by splitting the matrix, the expressive power can be increased under a fixed storage constraint. For example, if the matrix is divided into $m$ blocks, each with $c$ centroids, the entire matrix can express $c^{m}$ centroids while only storing $c\times m$ centroids. Without splitting, storing $c^{m}$ centroids would be required.
3.2 HETA Quantization
The proposed HETA Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness. The algorithm proceeds in four key steps, as shown in Algorithm 2:
@@ 3.2 HETA Quantization
Algorithm 1 HETA Quantization Algorithm
1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of HETA iterations.
@@
3.2.1 HETA Dequantization
The dequantization process, as described by the HETA Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$. This process involves several key steps, as shown in Algorithm 2:
@@ 3.2.1 HETA Dequantization
Algorithm 2 HETA Dequantization Algorithm
@@ 3.3 Optimizations
In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO). RO takes advantage of the fact that, after applying HETA, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE. On the other hand, BPO utilizes batch processing techniques to enhance processing speed. Since the codebook used in the HETA algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
@@ 3.3.1 Residual Optimization
Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the HETA algorithm. As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range. Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO). The RO algorithm, as depicted in Figure 1, begins with the quantization of the original matrix to produce a quantized matrix. This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix. The residual matrix typically exhibits a significantly reduced data range, making it easier to compress. This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or HETA.
@@ 3.3.2 Batch Processing Optimization
The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the HETA algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
@@ 3.3.3 Codebook Quantization Optimization (CQO)
The codebook occupies a significant portion of the space in the HETA algorithm. As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space. Therefore, compressing the codebook can significantly reduce the space used by the algorithm. Our optimization mainly employs the RTN method to compress the codebook.
4 Theory
@@ 5.1 Experiment setup
Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset. Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text. The results for the other datasets are provided in the appendix (Section A.1). Below, we provide a detailed description of each dataset.
@@
(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks. This dataset is particularly useful for evaluating the algorithm’s effectiveness in handling high-dimensional and dense data representations. (Jegou et al. (2010)) (Oliva & Torralba (2001))
(5) Wiki dataset: The Wiki dataset contains document embeddings derived from Wikipedia articles. This dataset tests the algorithm’s ability to compress and process large-scale text data efficiently.
@@
Metrics:
Compared algorithms:
Comparison of Algorithms and Parameter Selection
@@
Our HERA algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
$k_{s}\times d\times 32+n\times m\times[(k_{s}-1)\cdot\text{bitlength}]\leq\text{memory in bits}$
@@
Comparison Metrics
We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) for experimental evaluation. Let $x_{(i,j)}$ and $x_{(i,j)}^{\prime}$ denote the values before and after dequantization, respectively. The following metrics are used for comparison:
$\text{MAE}=\frac{1}{n\cdot d}\sum_{i,j}\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|$
$\text{MRE}=\frac{1}{n\cdot d}\sum_{i,j}\frac{\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|}{x_{(i,j)}}$
$\text{MSE}=\frac{1}{n\cdot d}\sum_{i,j}\left(X_{(i,j)}-X_{(i,j)}^{\prime}\right)^{2}$
We performed an optimization using 1-4 iterations, corresponding to our algorithms labeled as our1 through our4. In the quantification accuracy experiment, we tested the accuracy with different numbers of groups $m$ and various sizes of data matrices. In the sensitivity experiment, we evaluated the impact of different numbers of groups $m$ on the results while maintaining the same compression ratio.
@@
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  4-  4, Figures  7-  7, Figures  10-  10, Figures  13-  13 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1616, Figures  1919, and Figures  $2$, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).

### v3

# Quantum Entanglement Trees: Optimizing Quantized Matrix Quantization via Element Replacement and Residual Clustering

@@ Abstract
We propose Quantum Entanglement Trees (QET) to address the QEM problem by leveraging the local orderliness of matrix elements, involving iterative element swapping to form a locally ordered matrix. This matrix is then grouped and quantized by columns. To enhance QET, we introduce two optimizations: further quantizing residuals to reduce MSE, and using masking and batch processing to accelerate the algorithm.
Experimental results demonstrate that QET can effectively reduce MSE to 12.3% of the original value at the same compression ratio, outperforming the best baseline algorithms. Our contributions include the abstraction of the QEM problem, the design of the QET algorithm, and the proposal of two optimizations to improve accuracy and speed.
@@ 1 Introduction
There has been a growing literature on the matrix quantization. The first category of methods focuses on independently compressing the elements of a matrix for each specific scenario. Examples of such techniques include LLM.int8 (Ge et al. (2013)), Optimal Brain Damage (LeCun et al. (1989)), GPTQ (Frantar et al. (2022)), AWQ (Lin et al. (2023)), SmoothQuant (Xiao et al. (2023)), and OmniQuant (Shao et al. (2023)). The second category groups matrix elements by columns and then applies quantization to each group. Relevant works in this category include product quantization (PQ) (Jegou et al. (2010)), optimized product quantization (OPQ) (Ge et al. (2013)), locally optimized product quantization (LOPQ) (Kalantidis & Avrithis (2014)), and SqueezeLLM (Kim et al. (2023)). The first category of work can be summarized as utilizing outliers and the importance of elements in specific scenarios to improve the RTN (Round-To-Nearest)  [^1] RTN (Round-To-Nearest) is an algorithm that quantizes elements by rounding each value to its nearest representable level. (Gray & Neuhoff (1998)) algorithm. The second category focuses on enhancements to PQ algorithms. The primary drawback of the first category is that it does not consider the correlation between elements, instead independently quantizing and storing elements. Conversely, the second category fails to account for the relative order of elements.
@@
Using these ideas, we propose Quantum Entanglement Trees (QET) to address the QEM problem in matrix quantization. The core idea of QET is to leverage the local orderliness of matrix elements to optimize the QEM problem. The design involves swapping adjacent elements of the matrix to form a new, locally ordered matrix. To cover a broader range, we can perform further element swapping on the initially locally ordered matrix, similar to the approach of receptive field used in convolutional neural networks (LeCun et al. (1998)). This step can be repeated for multiple iterations. The newly ordered matrix is then grouped by columns, with each group being quantized.
Additionally, we propose two optimization algorithms based on the basic QET. First, the residuals between the original and quantized matrices can be further quantized to enhance the accuracy of the results. Second, since the iterative swapping operations can slow down the algorithm, we introduce optimization techniques based on masking and batch processing to accelerate the QET.
We evaluate QET on multiple datasets. Experimental results demonstrate that our method can effectively reduce the MSE to 12.3% of the original value at the same compression ratio, compared to the best baseline algorithm, which achieves an MSE of XXXX%.
@@
  • Designed an algorithm: We developed the Quantum Entanglement Trees (QET) algorithm, leveraging the concept of local orderliness to optimize the QEM problem.
  • Proposed two optimizations: We introduced the use of residuals to reduce MSE and employed batch processing techniques to accelerate QET algorithm.
@@ 3 METHODS
The core idea of Quantum Entanglement Trees (QET) is to leverage the local orderliness of matrix elements by rearranging adjacent elements to optimize the matrix quantization algorithm. Additionally, multiple iterations are used to expand the algorithm’s swapping field. We begin by describing the basic version of this idea and then propose two optimizations for the basic algorithm.
@@ 3.1 Basic algorithm
The steps and design of the QET algorithm are illustrated in Figure 2. The figure provides a conceptual overview of the Quantum Entanglement Trees (QET) algorithm.
On the left side of the figure, the process of rearranging matrix elements through multiple iterations in the QET is depicted. This iterative procedure involves comparing adjacent elements and ensuring that adjacent elements are not placed in the same matrix. This separation of adjacent elements into different matrices is conceptually similar to quantum entanglement, hence the name ”Quantum Entanglement Trees.” The QET algorithm performs multiple iterations, with each iteration comparing adjacent elements and splitting them based on their values. This approach is based on two observations. First, the local orderliness of the matrix can enhance the regularity of matrix elements, thereby improving compression efficiency. Second, multiple iterations allow the orderliness to cover a larger number of elements. For instance, adjacent elements after the first iteration might be non-adjacent in the original matrix. This iterative process, similar to that of convolutional neural networks, extends the ”receptive field” to cover more elements.
On the right side of the figure, the QET compression process is shown. After the elements have been rearranged, the algorithm groups the sub-matrices, followed by clustering to determine the centroids. These centroids are used to quantize the matrix. The final quantized matrix, codebook, and indicator maps are the outputs of the QET algorithm. This divide-and-conquer compression method is based on an observation: by splitting the matrix, the expressive power can be increased under a fixed storage constraint. For example, if the matrix is divided into $m$ blocks, each with $c$ centroids, the entire matrix can express $c^{m}$ centroids while only storing $c\times m$ centroids. Without splitting, storing $c^{m}$ centroids would be required.
3.2 QET Quantization
The proposed QET Algorithm is designed to optimize the quantization of a matrix $X\in\mathbb{R}^{n\times d}$ by leveraging local orderliness. The algorithm proceeds in four key steps, as shown in Algorithm 2:
@@ 3.2 QET Quantization
Algorithm 1 QET Quantization Algorithm
1:Input: Matrix $X\in\mathbb{R}^{n\times d}$, $m$ is the number of subspaces for quantization, $l$ is the number of QET iterations.
@@
3.2.1 QET Dequantization
The dequantization process, as described by the QET Dequantization Algorithm, aims to reconstruct the original matrix $X^{\prime}$ from its quantized version $X^{q}$. This process involves several key steps, as shown in Algorithm 2:
@@ 3.2.1 QET Dequantization
Algorithm 2 QET Dequantization Algorithm
@@ 3.3 Optimizations
In this section, we propose three optimizations: Residual Optimization (RO), Batch Processing Optimization (BPO), and Codebook Quantization Optimization (CQO). RO takes advantage of the fact that, after applying QET, the range of the data is relatively small, allowing for further optimization of the residual matrix and improvement of the MSE. On the other hand, BPO utilizes batch processing techniques to enhance processing speed. Since the codebook used in the QET algorithm occupies significant space, CQO further quantizes the codebook to improve space efficiency.
@@ 3.3.1 Residual Optimization
Residual Optimization (RO) is grounded in our observation that the data range of the quantized matrix is significantly reduced after applying the QET algorithm. As shown in Table 1, our experiments indicate that at a compression ratio of 16, the data range of the quantized matrix is reduced to 29.7% of the original matrix’s range. Moreover, with a compression ratio of 4, the data range of the quantized matrix is reduced to approximately 1.6% of the original matrix’s range.
Matrix quantization algorithms are more efficient at compressing matrix with a smaller data range; therefore, we propose Residual Optimization (RO). The RO algorithm, as depicted in Figure 2, begins with the quantization of the original matrix to produce a quantized matrix. This matrix is then dequantized to yield a dequantized matrix, which is subtracted from the original matrix to form a residual matrix. The residual matrix typically exhibits a significantly reduced data range, making it easier to compress. This residual matrix is then quantized using a chosen matrix quantization algorithm, such as RTN or QET.
@@ 3.3.2 Batch Processing Optimization
The key insight behind (Batch Processing Optimization) BPO is that by organizing and processing data in larger chunks or batches and enhancing the efficiency of the QET algorithm through batch processing and vectorization operations, the algorithm can more efficiently determine the relationships between elements, such as identifying smaller and larger parts or setting indicator maps.
@@ 3.3.3 Codebook Quantization Optimization (CQO)
The codebook occupies a significant portion of the space in the QET algorithm. As shown in Table 1, our experiments indicate that when the compression ratio is 16, the codebook occupies 62.5% of the total space, and when the compression ratio is 4, the codebook occupies 87.5% of the total space. Therefore, compressing the codebook can significantly reduce the space used by the algorithm. Our optimization mainly employs the RTN method to compress the codebook.
4 Theoretical Guarantees of QET
In this section, we will first derive the MSE of the PQ and QET algorithms, then theoretically prove that the MSE of our QET algorithm is lower than that of the PQ algorithm.
Theorem 1.
Suppose each element in the matrix is independently sampled from a normal distribution $x_{(i,j)}$ $\sim$ $\mathcal{N}(\mu,\sigma)$. After Quantizated by the PQ algorithm, the MSE is as follows:
$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$
$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
[…]
$C_{ij}\quad\text{residual matrix}$
$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
$\text{MSE}=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
$\text{MSE}=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (1)
In this section, this paper will theoretically prove that the MSE of our HERA algorithm is lower than that of the PQ algorithm.
Suppose each element in the matrix is independently sampled from a normal distribution $\mathcal{N}(\mu,\sigma)$. After compression by the PQ algorithm, the process for calculating MSE is as follows:
$A_{ij}\sim\mathcal{N}(\mu,\sigma^{2})\quad\text{independently and identically distributed}$
[…]
$C_{ij}\quad\text{residual matrix}$
$E[C_{ij}]=E[A_{ij}-B_{ij}]=\mu-\mu=0$
$MSE=\text{Var}[C_{ij}]=\text{Var}[A_{ij}]+\text{Var}[B_{ij}]-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\text{Cov}(A_{ij},B_{ij})$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2E\left[(A_{ij}-\mu)\left(\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}A_{ij}-\mu\right)\right]$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\left[\frac{1}{n_{k}}E\left[(A_{ij}-\mu)^{2}\right]+\frac{1}{n_{k}}\sum_{i\neq i^{\prime}}E\left[(A_{ij}-\mu)(A_{i^{\prime}j}-\mu)\right]\right]$
$\text{Since}\ A_{ij}\text{ and}\ A_{i^{\prime}j}\ (i\neq i^{\prime})\ \text{are independent}$
$=\sigma^{2}+\frac{\sigma^{2}}{n_{k}}-2\cdot\frac{\sigma^{2}}{n_{k}}$
$=\sigma^{2}-\frac{\sigma^{2}}{n_{k}}$ (2)
For the HERA algorithm proposed in this paper, the MSE obtained using the HERA algorithm on a matrix with elements following a $\mathcal{N}(\mu,\sigma^{2})$ distribution is equivalent to the MSE obtained using the PQ algorithm on a matrix with elements following a distribution with variance $0.682\sigma^{2}$. According to formula equation 2 equation 3, the MSE of HERA is smaller than that of the PQ . The following discussion in this paper will explain why this equivalence holds.
For a Gaussian distribution, after applying element pairing in the HERA algorithm, the cumulative distribution function $F_{Y}(y)$ of the elements on the right side of the matrix (the side storing the larger elements) satisfies:
$F_{Y}(y)=P(Y\leq y)=P(\max(X_{1},X_{2})\leq y)=P(X_{1}\leq y)\cdot P(X_{2}\leq y)$
$\text{Each}\ X_{i}\text{ follows a normal distribution}\ N\ (\mu,\sigma^{2})$
$F_{Y}(y)=\Phi\left(\frac{y-\mu}{\sigma}\right)\cdot\Phi\left(\frac{y-\mu}{\sigma}\right)=\left[\Phi\left(\frac{y-\mu}{\sigma}\right)\right]^{2}=\frac{1}{4}\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]^{2}$
Take the derivative of $\text{Take the derivative of }F_{Y}(y):$
$f_{Y}(y)=\left[1+\text{erf}\left(\frac{y-\mu}{\sigma\sqrt{2}}\right)\right]\cdot\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$
The mean and variance of $\text{The mean and variance of }Y\text{ are given by:}$:
[…]
Use numerical integration methods:
$E(Y)=\mu+0.564\sigma$
$Var(Y)=0.682\sigma^{2}$ (3)
Taking symmetry into account, the left side of the matrix also adheres to equation equation 3.
@@ 5.1 Experiment setup
Dataset: To evaluate the effectiveness of our proposed algorithm, we conducted experiments using one synthetic dataset and four real-world datasets: a synthetic normal distribution dataset, LLM weight dataset, KV cache dataset, Gist dataset, and Wiki dataset. Due to space constraints, we present the results for the synthetic normal distribution dataset and the LLM weight dataset in the main text. The results for the other datasets are provided in the appendix (Section LABEL:sec:appexponaddi). Below, we provide a detailed description of each dataset.
@@
(4) Gist dataset: The Gist dataset consists of feature vectors used for image recognition and retrieval tasks. GIST consists of the first 100k features extracted from the tiny image set (Torralba et al. (2008)). The dataset is also used in Jegou et al. (2010) and Oliva & Torralba (2001).
(5) Wiki dataset: The Wiki dataset (wik (2023)) contains document embeddings derived from Wikipedia articles. This dataset includes 256-dimensional vectors for a sample of 1 million Wikipedia entries.
@@
Metrics: We primarily measure the accuracy and time consumption of the algorithm. We use MAE (Mean Absolute Error), MRE (Mean Relative Error), and MSE (Mean Squared Error) as accuracy metrics. QTR (Quantization Time Ratio) and DTR (Dequantization Time Ratio) are used as time consumption metrics. Below, we introduce these metrics in detail.
Let $x_{(i,j)}$ denote the elements of the original matrix to be quantized, and $x_{(i,j)}^{\prime}$ denote the elements of the dequantized matrix.
$\text{MAE}=\frac{1}{n\cdot d}\sum_{i,j}\left|x_{(i,j)}-x_{(i,j)}^{\prime}\right|,\;\text{MSE}=\frac{1}{n\cdot d}\sum_{i,j}\left(x_{(i,j)}-x_{(i,j)}^{\prime}\right)^{2}$
For time efficiency, we define the following metrics:
$\text{QTR}=\frac{T_{\text{quant}}}{T_{\text{quant, baseline}}},\;\text{DTR}=\frac{T_{\text{dequant}}}{T_{\text{dequant, baseline}}}$
In these equations, $T_{\text{quant}}$ and $T_{\text{dequant}}$ represent the time taken by the proposed algorithm to quantize and dequantize the data, respectively. $T_{\text{quant, baseline}}$ and $T_{\text{dequant, baseline}}$ refer to the corresponding times taken by the baseline algorithm.
Comparative Algorithms: For the abstract Quantization Error Minimization (QEM) problem, our comparative algorithms fall into two main categories. The first category involves independently compressing the elements of a matrix for each specific scenario, which can be abstracted as the RTN (Round-To-Nearest) algorithm. The second category groups matrix elements by columns and then applies quantization to each group. Related algorithms in this category include PQ, OPQ, and LOPQ. Therefore, our comparative algorithms are as follows: RTN (Round-To-Nearest) (Gray & Neuhoff (1998)), PQ (Product Quantization) (Jegou et al. (2010)), OPQ (Optimized Product Quantization) (Ge et al. (2013)), and LOPQ (Locally Optimized Product Quantization) (Kalantidis & Avrithis (2014)).
Parameter Selection
@@
For Residual Optimization (RO), the large space required by RTN makes it infeasible for the algorithm to achieve high compression ratios. Therefore, in our experiments, we primarily use QET as the residual algorithm for RO.
Our QET algorithm involves a critical parameter, the number of centroids $k_{s}$. The selection of $k_{s}$ is governed by the following memory constraint:
@@
We measure MAE, MRE and MSE on different numbers of subspaces to show the accuracy of HERA.The results for MAE, MRE and MSE for different number of subspaces are shown in Figures  5-  5, Figures  8-  8, Figures  11-  11, Figures  14-  14 respectively. For 8-subspace setting, HERA can reduce the MSE to 70.4%, 49.7%, 35.1% and 12.3% in 1-4 iterations, respectively.
@@
We also conducted a parameter sensitivity test under different iterations, evaluating how the number of subspaces affects the algorithm’s performance. The results for 1 to 4 iterations are shown in Figures  1717, Figures  2020, and Figures  2323, respectively. Extensive experiments demonstrate that our algorithm reduces parameter sensitivity, making parameter selection more user-friendly (as indicated by the closer proximity of parameters in our algorithm’s curves).
6 Future Work
In future work, we plan to apply this algorithm to the quantization of large model parameters and the kv cache. This will allow us to evaluate the algorithm’s effectiveness in handling the challenges posed by large-scale models and real-time processing scenarios, further extending its applicability and performance in practical machine learning tasks.